Write the equation of the line that passes through (2, 2) and (6, 3) in standard form
Thank you

Answer by mananth(16946)

The slope of line passing through these points
y2-y1 / x2-x1
=(3-2) / (6-2)
=1/4 .
y= mx+b is the equation of the line.
plug any value from the two points.
2=1/4 * x+b
2=1/2 +b
2-1/2 = b
3/2 = b
..
so the equation of the line = y=1/4 (x) + 3/2